Use the labor-leisure model to analyze the effect on hours of work of a law requiring employers to pay “time-and-one-half” (that is, 1.5 times the straight-time wage) for hours worked over 40 per week. (A) Draw a budget constraint for someone earning $8.00 per hour who receives $100 a week in non- labor income. And total available hour is 100 per week excluding sleeping or going to church. What now happens to the budget constraint with the time-and-one-half requirement? (B) Given your answer to (A), draw an indifference curve so that, given the original budget constraint, the person is working 40 hours per week. What now happens to hours of work given the new budget constraint? Draw an indifference curve to show this phenomenon. Have hours of work increased or decreased as a result of the overtime law? Explain why. (C) In words, define the income and substitution effects. Next, use the graph to identify the income and substitution effects. Which one is stronger?


1
Expert's answer
2021-06-18T11:57:23-0400

Labour -Leisure diagram

L represents Leisure and l show the labour

"l+L=\bar{L}" (Total available hours)



Assuming 8 hours a day and 6 days a week

"\bar{L}=8*6=48 hours"

Budget line

"C+40 L = 40 \bar{L}"

where C is the consumption

"C+40 L = 40 *48 = 1920"

New budget line

for "0<L<40 or 48 > L> 40"

"C+20 L = 40* 20"

Beyond 40

"C+30 L =40*20+30*8"

"C+30 L =1040"






From available points on A below B were available before the person will not choose those points on A below B

The optimal choice could be at L = 8 as well as L = 40

The entire change will be a substitution effect. This is because, in substitution effect, the choose is found on the adjusted budget line

Here adjusted budget line is the new budget line thus the entire effect in substitution effect.